OpenStudy (anonymous):

Suppose G(x) = f(u(x)) and each of f and u are differentiable functions. If f (1) = 2, f (3) = −1, f '(1) = −1, f '(3) = 5, u(2) = 3, u '(2) = 4 , then find G'(2)

OpenStudy (anonymous):

G'(x)=f'(u(x))*u'(x) G'(2)=f'(u(2))*u'(2)=f'(3)*4=4*5=20

OpenStudy (anonymous):

that's what I got! sweet!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

got one more coming up, I got -3 for this one

OpenStudy (anonymous):

throw it... i'll try to help...

OpenStudy (anonymous):

-3 for the next one

OpenStudy (anonymous):

thanks a million!

OpenStudy (anonymous):

it's up, looks the same but isn;t